How to find the area of a parallelogram with the given vertices. P(1,3) Q(3,3) R(7,8) S(9,8). Please explain.
Answer by solver91311(24713)



The difference in the x-coordinates of either of the bases is 2; 3 - 1 = 2 and 9 - 7 = 2, hence the measure of the base is 2. The difference in the y-coordinates is 8 - 3 = 5, so the altitude is 5. Use the formula for the area of a parallelogram.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism